OpenStudy (marco26):

Hi! Need help regarding calculus. The region bounded by the curve y=x^2 +1 and the line y=x+3. Find the area of the region using horizontal strip. Find also the volume of the region if the axis of revolution is y=8. Thanks!

OpenStudy (anonymous):

www.most.gov.mm/techuni/media/EM_02011_chap456.pdf

OpenStudy (dumbcow):

@rohangrr, please let people know what the link is you are posting if it is a huge pdf document (41 pages seems excessive)

OpenStudy (dumbcow):

 \[A = \int\limits_{-1}^{2}(x+3)-(x^{2}+1) dx\]

OpenStudy (dumbcow):

For volume, you are summing up the areas of the cross-sections each cross-section is a ring with inner radius and outer radius inner = 8-(x+3) = 5-x outer = 8-(x^2+1) = 7-x^2 \[V =\pi \int\limits_{-1}^{2}(7-x^{2})^{2}-(5-x)^{2} dx\]

OpenStudy (marco26):

@dumbcow Thanks for the fast reply. But isn't that for the vertical strip for the area? .

OpenStudy (dumbcow):

oh yeah, sorry didn't think about that but since the functions are in terms of x its easier to integrate wrt x

OpenStudy (marco26):

I was able to compute that for the vertical strip. I think the answer was 9/2. But for some reason I'm having a hard time for the horizontal strip. Must have 2 regions, I guess. Unfortunately, no luck, I'm having diff. answer with horizontal strip.

OpenStudy (dumbcow):

well because its a parabola, it makes it difficult for region: 2<y<5 \[A = \int\limits_{2}^{5}\sqrt{y-1} - (y-3) dy\] for region: 1<y<2 \[A = 2\int\limits_{1}^{2}\sqrt{y-1} dy\]

OpenStudy (marco26):

Got it! Thanks